Tool Manufacturing has an expected EBIT of $71,000 in perpetuity and a tax rate of 23 percent. The company has $127,000 in outstanding debt at an interest rate of 6.8 percent and its unlevered cost of capital is 14 percent.

  

What is the value of the company according to MM Proposition I with taxes? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1
Value of unlevered firm = EBIT*(1-tax rate)/unlevered cost of equity
=71000*(1-0.23)/0.14
=390500
Value of levered firm = Value of unlevered firm + debt*tax rate
=390500+127000*0.23
=419710